

********************************************************************
UARS WEEKLY STATUS REPORT - 31 AUG 2001
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM has completed its daily solar scans this past week during
periods of predicted UARS telemetry.
The CDHF continues to reprocess SUSIM data and we are optimistic that
the job will be completed before the termination of reprocessing
operations.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
PEM is saddened by the shutdown of UARS and appreciates all of the
efforts of GSFC in the past and present of UARS.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha). Currently, PEM has requested GSFC assistance with compiling for
the DEC Alpha since the GSFC executables run on the Alpha and SwRI's compiled
versions do not.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in planning for the thruster
calibration burn which will be conducted on
6 September. This calibration burn is required so
that it will be possible for FDF to plan future
manuevers in case it becomes necessary to move
UARS to a lower altitude.
On 30 August 2001 the MPG participated in the weekly UARS
Technical Team Meeting. This series of meetings is being
held to discuss contingencies since it appears that FY 2002
funds will be insufficient to continue collection of
science data by the UARS instruments.
The instrument operations schedule for the period
11 September through 24 September was distributed on
26 August 2001. Current plans call for all UARS instruments
to be configured to a standby or off mode after
24 September. (Since the ACRIM II instrument uses very
little power and is quite sensitive to on/off commands,
the ACRIM instrument will remain turned on.)
Processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 18 August 2001 (Orbit 54316, GMT Day 230) through 24 August
2001 (Orbit 54420, GMT Day 236).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 230/0033 54316 Power Off
The beta angle ranged between the angles of 9.7 degrees to 31.3 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 46 degrees on 25 August 2001 (DOY 237).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.9 to 2.2 degrees C. Additional battery performance data for each
battery on 10 August 2001 and 17 August 2001 are provided in Appendix D.
The clock error ranged from Ð10.4 to 10.6- msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 16
¥ WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ WINDII = 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/235 show the observatory data loss to be 303 hours, 08
minutes, 57.696 seconds. This shows a loss of 02 minutes and 17.440 seconds
since the last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
The increase of 02 minutes and 17.440 seconds of data loss occurred over three
(3) days:
· 49.728 seconds on Day 230 due to a Data Hit
· 33.872 seconds on Day 231 due to a Data Hit
· 17.888 seconds on Day 233 due to a Data Hit
· 35.952 seconds on Day 235 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support and plan for the possible decommissioning of the
UARS spacecraft. A 3-minute calibration burn has been scheduled for 30 August
2001. The time of the burn will be set this Friday, but it most likely occur
during the first orbit after 1300 GMT. This burn should give the FOT and FDF a
good handle on the duty cycle and efficiency of the thrusters. All issues with
FDF product deliveries have been worked out and next Tuesday products should
include pre and predicted post burn data.
MEETINGS
The UARS FOT attended a UARS Flight Software CCB Meeting on Monday, August 20,
2001.
The UARS FOT attended a UARS FDF Meeting on Tuesday, August 21, 2001.
The UARS FOT attended a UARS Battery Meeting on Tuesday, August 21, 2001.
The UARS FOT attended a UARS SERB Meeting on Wednesday, August 22, 2001.
The UARS FOT attended a UARS Technical Team Meeting on Thursday, August 23,
2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 11 September 2001 (DOY 254). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 August 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
54316 230/0033 01-524 NBTR NOOPS D-230
54316 230/0033 01-525 WINDII Power OFF
54318 230/0331 01-524 NBTR NBTR PLAYBACK & RTS-58
54319 230/0517 01-524 NBTR NOOPS D-230
54326 230/1711 01-524 NBTR NOOP D-230
54357 232/1855 01-526 OBC Clk Adj 0.00 to 20.6 msec/day
54372 233/1755 01-527 OBC BANK DUMPS
54377 234/0250 01-529 OBC/CLOCK Clk Adj 20.6 to 0.00 msec/day
54380 234/0734 01-528 NBTR NBTR PBK & RTS 58 ACTIVATE
54391 235/0042 01-530 MLS EXCEPTION CLEAR
54392 235/0231 01-530 MLS EXCEPTION CLEAR
54393 235/0336 01-530 MLS EXCEPTION CLEAR
54393 235/0412 01-530 MLS EXCEPTION CLEAR
54394 235/0516 01-530 MLS EXCEPTION CLEAR (twice)
54398 235/1217 01-530 MLS EXCEPTION CLEAR
54404 235/1757 01-530 MLS EXCEPTION CLEAR
54417 236/1758 01-531 OBC Clk Adj fr 0.00 to 20.6 msec/day
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 August 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 August 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
54357 232/1855 Clk Adj fr 0.00 to 20.6 msec/day
54377 234/0236 Clk Adj fr 20.6 to 0.00 msec/day
54417 236/1807 Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 24 August 2001
17 August 2001 - GMT Day 229 Beta = 31.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-95.2 +5.47/+3.23 +28.0/-16.8 24.5 1.020 19.3
3 +22.4/-50.4 +4.51/+2.60 +27.2/-16.8 24.5 1.020 18.9
24 August 2001 - GMT Day 236 Beta = 45.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-123.2 +5.15/+4.19 +26.0/-14.4 25.4 1.030 19.7
3 +22.4/-50.4 +4.51/+3.23 +25.6/-14.4 25.4 1.020 19.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 August 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-051 WINDII 54316 230/0033 N/A WINDII non-responsive
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-051 WINDII 54316 230/0033 N/A WINDII non-responsive
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-045 FOT see list August 01 14170 Generic Late Acquisition
SEVNTFW01/33